OpenStudy (kingj):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 34 terms?

OpenStudy (mertsj):

\[\S _{n}=n(\frac{a _{1}+a _{n}}{2})\]

OpenStudy (mertsj):

What is asubn???

OpenStudy (mertsj):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (mertsj):

\[a _{34}=3+(33)(6)=201\]

OpenStudy (mertsj):

\[\S _{n}=34(\frac{3+201}{2})=3468\]